OpenStudy (anonymous):

Help will give Medal Which of the following equations would graph a line perpendicular to y = 3/4x + 5? A. y =3/4 x – 5 B. y = -3/4x + 5 C. y = 4/3x + 5 D. y = -4/3x + 5

OpenStudy (anonymous):

a perpendicular line will have a negative reciprocal slope. All that means is " flip " the slope and change the sign. Example : 5/6...negative reciprocal is -6/5 (see how I flipped the slope and changed the sign). slope 5/6 and slope -6/5 will be perpendicular. So what will be the negative reciprocal of the slope given in y = 3/4x + 5 ?

OpenStudy (anonymous):

Just realized my previous answers were of no help...! Kelliegirl33 is absolutely right here.

OpenStudy (anonymous):

Would it be C?

OpenStudy (anonymous):

not C. What is the slope in y = 3/4x + 5 ?

OpenStudy (anonymous):

in y = mx + b form, the slope is in the m position. So what is the slope in y = 3/4x + 5 ?

OpenStudy (anonymous):

wait i think i understand would it be D? because you flip 3/4 which would be 4/3 then it would be -4/3? (im not very good at math)

OpenStudy (anonymous):

that is correct....the negative reciprocal can be found by " flipping " and changing the sign. So negative reciprocal of 3/4 is -4/3....so your perpendicular equation is : y = -4/3x + 5

OpenStudy (anonymous):

Thank you! that makes so much more sense now!

OpenStudy (anonymous):

glad to help :)

OpenStudy (anonymous):

remember this...a parallel line will have the same slope, but perpendicular lines will have negative reciprocal slopes

OpenStudy (anonymous):

ok i will Thanks again!

OpenStudy (anonymous):

sure thing :)